\begin{thebibliography}{10}

\bibitem{IHF05}
 T. Iwasaki, S. Hara, A. Fradkov.  Time domain interpretations of frequency
domain inequalities on (semi)finite ranges. Systems \& Control
Letters, 2005. Vol 54 (7), pp 681-691.

%\bibitem{doyle:90book}
%J.C. Doyle, B.~Francis, and A.~Tannenbaum.
%\newblock {\em Feedback Control Theory}.
%\newblock Macmillan Publishing Co., 1990.

%\bibitem{green:95book}
%M.~Green and D.J.N. Limebeer.
%\newblock {\em Linear Robust Control}.
%\newblock Prentice Hall, New Jersey, 1995.


\bibitem{iwasaki:kyp}
T.~Iwasaki, G.~Meinsma, and M.~Fu.
\newblock Generalized {$S$}-procedure and finite frequency {KYP} lemma.
\newblock {\em Mathematical Problems in Engineering}, 6:305--320, 2000.

\bibitem{iwasaki:dyn}
T.~Iwasaki, S.~Hara, and H.~Yamauchi.
\newblock Dynamical system design from a control perspective: {F}inite
  frequency positive-realness approach.
\newblock {\em IEEE Trans. Auto. Contr.}, 48(8):1337--1354, August 2003.

\bibitem{iwasaki:gkyp}
T.~Iwasaki and S.~Hara.
\newblock Generalized {KYP} lemma: {U}nified frequency domain inequalities
with design applications.
\newblock {\em IEEE Trans. Auto. Contr.}, Vol.50, No.1, January 2005.

\bibitem{willems:72II}
J.~C. Willems.
\newblock Dissipative dynamical systems part {II}: {L}inear systems with
  quadratic supply rates.
\newblock {\em Archive of Rational Mechanics and Analysis}, 45(5):352--393, 1972.

\bibitem{yak73}
Yakubovich V.A.  Minimization of quadratic functionals under
quadratic constraints
   and the necessity of frequency condition in the quadratic criterion for
   absolute  stability of nonlinear control systems.
   Soviet Math.Dokl., 1973, v.14,  pp.593-597.
   (Translated from Doklady AN SSSR, 1973, v.209, pp. 1039-1042).

\bibitem{yakfr73}
Fradkov A.L.,Yakubovich V.A.  Necessary and sufficient conditions
for
   absolute  stability in classes of systems with two integral quadratic
   constraints.-Soviet Math.Dokl., 1973, v.14, N6, pp.1812-1815.
   (Translated from Doklady AN SSSR,1973, v.213, N 4).

%\bibitem{anderson:67}
%B.~D.~O. Anderson.
%\newblock A system theory criterion for positive real matrices.
%\newblock {\em SIAM J. Contr.}, 5:171--182, 1967.

%\bibitem{rantzer:kyp}
%A.~Rantzer.
%\newblock On the {K}alman-{Y}akubovich-{P}opov lemma.
%\newblock {\em Sys. Contr. Lett.}, 28(1):7--10, 1996.

%\bibitem{willems:71}
%J.~C. Willems.
%\newblock Least squares stationary optimal control and the algebraic {R}iccati
%  equation.
%\newblock {\em IEEE Trans. Auto. Contr.}, 16:621--634, 1971.

\bibitem{yakubovich:71}
V.~A. Yakubovich.
\newblock {S}-procedure in nonlinear control theory.
\newblock {\em Vestnik Leningrad Univ.}, 1:62--77, 1971.

\bibitem{F73}
A.L. Fradkov.
\newblock Duality theorems for certain nonconvex extremal problems.
\newblock {\em Siberian Mathematical Journal}, 14(2):357--383, March-April
  1973.
%\newblock (English translation).

\bibitem{Yak92}
V.~A. Yakubovich.
\newblock Nonconvex optimization problem: the infinite-horizon linear-quadratic
  control problem with quadratic constraints.
\newblock {\em Sys. Contr. Lett.}, 19:13--22, 1992.

\bibitem{megretski:93}
A.~Megretski and S.~Treil.
\newblock Power distribution inequalities in optimization and robustness of
  uncertain systems.
\newblock {\em J. Math. Syst. Estim. Contr.}, 3(3):301--319, 1993.

\bibitem{MatveevYak93}
Matveev A.S., Yakubovich V.A. Nonconvex problems of global
optimization, St.Petersburg Math. J., 1993,  pp.1217--1243.

\bibitem{Matveev98}
Matveev A.S., Spectral approach to duality in nonconvex global
optimization, SIAM J. Control and Optimization, vol.36 (1998),
No.1, pp.336-378.


\end{thebibliography}

\documentclass{article}
\usepackage{fleqn,times,cite}
\usepackage[intlimits]{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{hhline}

\twocolumn

\newtheorem{defn}{Definition}
\newtheorem{theorem}{\sf\bfseries Theorem}
\newtheorem{lemma}{\sf\bfseries Lemma}

\newtheorem{coro}{\sf\bfseries Corollary}
\def\IR{{\rm I\!R}}
\def\IC{{{\rm C}\!\!\!{\sf I}~}}
\def\QED{~\rule[-1pt]{5pt}{5pt}\par}

\newenvironment{mat}{\left[\begin{array}}{\end{array}\right]}

\renewcommand{\t}{^{\mbox{\tiny\sf T}}}
\newcommand{\dt}{{\mbox{\rm\em dt}}}

\newcommand{\hs}{\hspace{4mm}}
\newcommand{\hsm}{\hspace{-1.4mm}}
\newcommand{\he}{{\rm He}}

\newcommand{\bfOme}{{\bf\Omega}}

\newcommand{\Ltwo}{{{\cal L}_2[0,\infty)}}

\renewcommand{\j}{{\jmath}}

\newcommand{\inte}{{\rm Int}}

\newcommand{\mylabel}[1]{\label{#1}}
\newcommand{\mR}{\mathbb R}
\newcommand{\mZ}{\mathbb Z}

\title{\large\bf CONIC S-PROCEDURE AND CONSTRAINED DISSIPATIVITY
\\
\thanks{
 The work is supported in part by the Russian Foundation Foundation for Basic Research
(project 05-01-00869) and CNRS-RAS cooperative research program
(project 04-16394)} }

\author{~~~ Alexander L. Fradkov\thanks{
A. L. Fradkov is with the Institute for Problems of Mechanical
Engineering, Control of Complex Systems Laboratory, V.O., Bolshoy, 61,
St. Petersburg, 199178, Russia, {\rm alf@control.ipme.ru}}
}

\begin{document}
\date{}
\maketitle

\begin{abstract}
A new version of classical S-procedure in system theory is
proposed based on duality in the space of positive definite
matrices and introduction of matrix Lagrange multipliers. A new
proof and extension of the recent results \cite{IHF05}
%of T.Iwasaki, S. Hara, A. Fradkov ( Systems & Control Letters, 2005. Vol 54 (7), pp 681-691)
concerning equivalence between frequency domain inequality on
finite frequency range and constrained dissipativity property for
linear systems is given. The results of this paper  extend
S-procedure to allow for analysis and design of robust systems
with matrix inequalities constraints.
\end{abstract}

{\bf Keywords.} Kalman-Yakubovich-Popov (KYP) lemma, S-procedure,
frequency domain inequality, linear matrix inequality,
dissipativity.

\section{Introduction} \mylabel{sec:intro}

Recently a number of new tools for systems analysis and design
 related to frequency domain inequalities (FDI) over a finite frequency range
(so called {\it Generalized KYP-lemma}) have been developed
\cite{iwasaki:kyp,iwasaki:dyn,iwasaki:gkyp}. It follows from the
results of \cite{iwasaki:kyp,iwasaki:dyn,iwasaki:gkyp} that
fulfillness of a standard FDI in a finite frequency range is
equivalent to validity of some nonclassical linear matrix
inequalities (LMI) for a pair of matrices $P,Q$ replacing
inequalities for a single matrix $P$ appearing in the classical
KYP-lemma. It was shown in \cite{IHF05} that FDI, in turn, are
equivalent to some time-domain inequality (TDI, dissipation
inequality \cite{willems:72II}), valid only over a part of the
system trajectories, determined by an additional integral matrix
inequality (restricted or constrained dissipativity \cite{IHF05}).
Thus, a complete extension of the classical KYP-results on
equivalence between FDI, TDI and LMI to the ''finite-frequency"
case was obtained. Note that the proof of equivalence between FDI
and TDI in \cite{IHF05} goes along the
 lines of the necessity proof for the frequency-domain absolute stability criterion
\cite{yak73,yakfr73}.

In this paper a new proof of the result of \cite{IHF05} is
provided based on the losslessness result for a new version of the
classical S-procedure \cite{yakubovich:71}. A new version of the
S-procedure, also included in the paper, deals with constraints in
LMI form, or more generally, conic inequalities in linear spaces.

In the next section a new S-procedure results are presented. In Section 3 they are
 applied to the proof of equivalence between TDI and LMI.

We use the following notation. The set of square integrable
functions on $[0,\infty)$ is denoted by $\Ltwo $, $M^\dagger$,
where $M$ is a matrix, stands for  its transposition and complex
conjugate of all elements. For a square matrix $M$, its Hermitian
part is defined by $\he(M):=(M+M^\dagger)/2$.
%For a Hermitian matrix $M$, its
%maximum eigenvalue is denoted by $\lambda_{\max}(M)$. The
%operators $\Re(\cdot)$ and $\Im(\cdot)$ take the real and
%imaginary parts of the arguments, respectively. The set of
%positive integers up to $n$ is denoted by $\I_n$.
 The interior of a set $\bfOme$ is denoted by $\inte~\bfOme$.

\section{Conic S-procedure}
Let $X$, $Y_1$, $\dots, Y_m$ be linear topological spaces, $G_j\!:X\to\! Y_j$,
$j=1,\dots,m$ be continuous mappings.

Let for any $j=1,\dots,m$ a convex cone $K_j\subset Y_j$ be given
defining inequality $G_j(x)\ge 0$ for $x\in X$ as inclusion
$G_j(x)\in K_j$. Let $Y^*_j$ denote an dual space to $Y_j$, i.~e.
a linear space of linear continuous functionals $y^*_j$ on $Y_j$
and $K^*_j\!\subset\! Y^*_j$ denote an dual cone to $K_j$, i.~e.
$K^*_j\!=\!\big\{y_j^*\in Y_j^*: \big<y_j^*,y_j\big>\ge 0~ \forall
y_j\in K_j\big\}$, where $\big<y_j^*,y_j\big>$ is the value of the
functional $y_j^*$ at the element $y_j$.

Obviously, if $Y=\mR^1\times Y_1\times \dots \times Y_m$, then
$Y^*= \mR^1\times Y^*_1\times\dots\times Y^*_m$ is the set of all
corteges $\big(y_0^*,y_1^*,\dots,y_m^*\big)$, where $y_0^*\in
\mR^1$, $y_j^*$ is a linear functional from $Y_j^*$.

Consider the following two relations for the mappings $F_0$,
$G_1$, \dots, $G_m$.
\begin{itemize}
\item[({\it A})] $F(x)\ge 0$ for $x\in X$, $G_j(x)\in K_j$,
$j=1,\dots,m$; \item[({\it B})]$\exists\tau_0\ge 0$, $\tau_j\in
K^*_j$:
$\tau_0F(x)-\sum\limits_{j=1}^{m}{\big<\tau_j,G_j(x)\big>\ge 0} ~
 \forall x\in X$.
\end{itemize}
Obviously, validity of ({\it B}) with $\tau_0 > 0$ implies ({\it
A}). Indeed, if $x\in X$ satisfies inequalities $G_j(x)\in K_j$,
$j=1,\dots,m$, then it follows from ({\it B}) that $\tau_0F(x)\ge
0$, since $\big<\tau_i,G_i(x)\ge 0\big>$ for $j=1,\dots,m$. The
opposite statement is not true even in the case of scalar
constraints $Y_j=\mR^1$, $j=1,\dots,m$, corresponding to the
classical $S$-procedure \cite{yakubovich:71}.

Similarly to the classical case we will say that $S$-procedure
with conic constraints $G_j(x)\ge 0$ is lossless, if ({\it B})
with $\tau_0 > 0$ implies ({\it A}).

It is well known \cite{F73} that losslessness of the classical
$S$-procedure is equivalent to the duality theorem in the
corresponding optimization problem. However, the problem is, in
general, nonconvex and only a few classes of functionals $F$,
$G_1$, \dots, $G_m$ are known to possess the losslessness
property.

For example, classical $S$-procedure is lossless, if $m=1$ and
$F$, $G_1$ are quadratic forms on real or complex linear space
$X$. It is also lossless, if $m=2$ and $F$, $G_1$, $G_2$ are
quadratic (Hermitian) forms on the complex linear space $X$.
However, classical $S$-procedure for quadratic forms is, in
general, lossy for $m\ge 2$ in real case and for $m\ge 3$ in
complex case \cite{F73}. A.~Megretski and S.~Treil proved in 1990
\cite{megretski:93} that the classical $S$-procedure is lossless
for all $m\ge 1$, if $F$, $G_1$, \dots, $G_m$ are integral
quadratic forms on ${\cal L}_2(0, \infty)$. V.~Yakubovich extended
this result to a more broad class of quadratic functionals,
forming the so-called {\it S-system} \cite{Yak92}.

Below an extension of the results of \cite{Yak92} to the case of
the $S$-procedure with conic constraints is formulated. Note that
the general formulation of the $S$-procedure with conic
constraints was presented, e.g. in \cite{MatveevYak93,Matveev98},

\begin{theorem}\label{theorem:1}.
Let $n_0=1$ and $\bar K$ is  the closure of the cone $K$ generated
by the set

${\cal F} (X)=\{(F_0(x), F_1(x),\dots, F_m(x)):~x\in X\}. $

If the cone $\bar K$ is convex, then the S-procedure with conic
constraints is  lossless.

If, in addition, constraints $G_j(x)\in K_j$ are regular, namely
$\exists x_0:~G_j(x_0)\in \inte~K_j$, then one can choose
$\tau_0=1$ in (B).
 \end{theorem}

{\it Proof}. Condition (A) implies that $F(x)\ge 0$ for $G_j(x)\in
\inte K_j$, i.e. intersection of the set ${\cal F}(X)$ and the
open cone $D=\{(-y_0, y_1, \dots, y_m): y_0>0, y_j\in \inte K_j,
j=1,\dots,m\}$ is empty: $D\bigcap {\cal F}(X)=\phi$. Therefore,
$D\bigcap\bar K=\phi$. Applying separation theorem for cones, we
obtain that there exists vector
$\tau^*=(\tau_0^*,\tau_1^*,\dots,\tau_m^*)\in Y^*$ such that
$<\tau_0^*,F(x)>+\sum_{j=1}^m<\tau_j^*,G(x)>\ge 0$ for all $x\in
X$ and $<\tau^*,y><0$ for all $y\in D$, i.e.
$<\tau_0^*,y_0>+\sum_{j=1}^m<\tau_j^*,y_j><0$. For any $j=1,\dots
m$ pick up $y_j\in \inte K_j$ and choose sequences
$y_{0_k}\rightarrow 0,~y_{s_k}\rightarrow 0$ as
$k\rightarrow\infty$, such that $y_{0_k}>0, y_{s_k}\in K_s, s\neq
j$. If $k\rightarrow\infty$, then we obtain $<\tau_j^*,y_j>\le 0$,
i.e. $-\tau_j^*\in K_j^*$. The first part of the theorem is
proved.

Taking $x=x_0$ from regularity condition and $y_0\neq 0$ yields
$\tau_0^*,y_0> 0$, i.e. $\tau_0^*>0$. Dividing the inequality (B)
by $\tau_0$, we arrive at the second statement of the theorem. End
of the proof.

 Our next step is to extend the definition of
$S$-system \cite{Yak92} to the case of conic constraints.

\begin{defn}. Let $F_j$, $j=0$, $1$, \dots, $m$ be mappings from
a Hilbert space $\mZ$ to spaces of self-adjoint operators over
corresponding Euclidean space $\mR^{n_j}$, such that $F_j: \mZ\to
SR(n_j\times n_j)$.

We say that $F_0$, $F_1$, \dots, $F_m$ form a S-system if there
exists a subspace $\mZ_0$ and a sequence of linear bounded
operators $T_k: \mZ\to\mZ$, $k=1$, $2$, \dots ~such that
\begin{itemize}
\item[({\it i})] $<T_kz_1,z_2>\to 0$ as $k\to\infty$ for all
$z_1$, $z_2\in \mZ$; \item[({\it ii})] $\mZ_0$ is invariant for
$T_k$ for all $k=1$, $2$, \dots; \item[({\it iii})] $F_j(T_kz)\to
F_j(z)$ as $k\to\infty$ for all $j=0$, $1$, \dots m, $z\in\mZ_0$.
\end{itemize}
\end{defn}

\begin{lemma} Let $F_j$, $j=0$, $1$, \dots, $m$ form $S$-system.
Define the map ${\cal F}:
\mZ\to\prod\limits_{j=0}^{n}{\mR^{n_j\times n_j}}$ by means of the
relation

\begin{align*}{\cal F}(z)&=F_0(z)\otimes F_1(z)\otimes
\dots \otimes F_m(z)\\&\in \mR^{n_0\times
n_0}\otimes\mR^{n_1\times n_1}\otimes\dots\otimes \mR^{n_m\times
n_m}.\end{align*}

Then the closure of the image ${\cal F}(\mZ)$ is a convex set in
$\mR^{n_0^2+n_1^2+\dots +n_m^2}$.\label{lemma:1}
\end{lemma}

In the special case $n_j=1$, $j=1, \dots, m$, Lemma \ref{lemma:1}
coincides with Lemma 1 of the paper \cite{Yak92} and it is proved
similarly to the Lemma 1 of \cite{Yak92}.

{\bf Example 1.} An important series of examples for $S$-systems
is provided by finite family of integral quadratic operators on
the Hilbert space ${\cal L}_2[0,\infty)$ of square integrable
functions with values $z(t)\in \mR^{n_j}$. The mappings are
defined for any $z\in {\cal L}_2[0,\infty)$ as follows:
\begin{equation}
\label{f1} F_j(z)=\he~\int\limits_0^\infty{F_j^{'}
z(t)z^{\dagger}(t)F_j^{''\dagger}\,dt,}
\end{equation}
where $F_j^{'},~F_j^{''}$ are $n_j\times n_j^{'}$ symmetric
matrices.
 In this case the family of the operators $T_k$ can be chosen
as time shifts: $T_k(z)(t)=z(t+k)$, while the subspace $\mZ_0$ can
be chosen as the set of functions with zero initial conditions:
\begin{align*}
\mZ_0=\big\{z(\cdot): z(\cdot)\in {\cal L}_2(0,\infty),
z_2(0)=0\big\}.
\end{align*}

The proof of the S-system property for Example 1 is again similar
to \cite{Yak92}. Note that the cone of positive semidefinite
matrices is selfdual. Therefore S-procedure with conic constraints
determined by functions (\ref{f1}) deals with positive
semidefinite matrix Lagrange multipliers.

Properties of the S-procedure in general case are given by the
following theorem.

\begin{theorem}\label{theorem:2}. $S$-procedure with conic
constraints is lossless for any family of self-adjoint operators
$F_0$, $F_1$, \dots, $F_m$ forming an $S$-system. \end{theorem}

Proof follows immediately from Theorem \ref{theorem:1} and Lemma
\ref{lemma:1}. The result can be extended to the case of equality
constraints and to the case of the  so called {\it generalized
$S$-procedure} introduced in \cite{iwasaki:kyp}.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Constrained dissipativity}

In this section,  we first present a special case of the
generalized KYP lemma \cite{iwasaki:gkyp},  characterizing FDIs in
the continuous-time setting. Let complex matrices $A$, $B$, $\Pi$,
and real scalars $\varpi_1$, $\varpi_2$ be given.
%Let $\tau$ be $+1$ or $-1$, and
Define
\begin{equation} \mylabel{Ome}
\bfOme:=\{~\omega\in\IR~|~(\omega-\varpi_1)(\omega-\varpi_2)\leq0~\}.
\end{equation} (We may assume $\omega_2>0$ without loss of
generality).

\begin{theorem} \mylabel{thm:ckyp} \cite{iwasaki:gkyp}.
Suppose $\Pi$ is Hermitian matrix, pair $(A,B)$ is controllable,
and $\bfOme$ has a nonempty interior. Then the following
statements are equivalent.
\begin{itemize}
\item[(i)] The frequency domain inequality
\begin{equation} \mylabel{fdic}
\begin{mat}{c} (j\omega I-A)^{-1}B \\ I \end{mat}^*\Pi
\begin{mat}{c} (j\omega I-A)^{-1}B \\ I \end{mat}\leq0
\end{equation}
holds for all $\omega\in\bfOme$ such that $\det(j\omega I-A)\neq0$.
\item[(ii)] There exist Hermitian matrices $P$ and $Q$ such that \\
$Q\geq0$ and the linear matrix inequality
\begin{equation} \label{lmic}
\begin{mat}{cc} \hsm A & \hsm B \hsm \\ \hsm I & \hsm 0 \hsm \end{mat}^*
\begin{mat}{cc} -Q & \hsm P+j\varpi_oQ \hsm \\
\hsm P-j\varpi_oQ & \hsm -\varpi_1\varpi_2Q \hsm \end{mat}
\begin{mat}{cc} \hsm A & \hsm B \hsm \\ \hsm I & \hsm 0 \hsm \end{mat}
+\Pi\leq0
\end{equation}
holds, where $\varpi_o:=(\varpi_1+\varpi_2)/2$.
\end{itemize}
\end{theorem}

%We mention that similar results have also been obtained in
%\cite{iwasaki:kyp,iwasaki:dyn}.
Choosing the parameters
$\varpi_1=\varpi_2=0$ and $\tau=-1$, the set $\bfOme$ becomes the
entire real numbers, and thus statement (i) becomes the FDI for
all frequencies. In this case, the term associated with $Q$ in the
LMI (\ref{lmic}) becomes positive semidefinite, and hence the best
choice of $Q$ for satisfaction of (\ref{lmic}) is $Q=0$. The resulting
LMI with variable $P$ is exactly the same as the one in
%\cite{rantzer:kyp}, and thus Theorem~\ref{thm:ckyp} reduces to
the standard KYP lemma.

The following result extends the result of \cite{IHF05}. It
provides an equivalence between FDI and time domain dissipation
inequality over a restricted class of input signals.

\begin{theorem} \mylabel{thm:tdic}.
Let complex matrices $A$, $B$, $\Pi$, and real scalars $\varpi_1$,
$\varpi_2$ be given and $\bfOme$ be defined by (\ref{Ome}).
Consider the system
\begin{equation} \mylabel{sysc}
\dot x(t)=Ax(t)+Bu(t), \hs t\in[0,\infty),
%\hs x(0)=0
\end{equation}
where $x(t)\in\IC^n$ is the state and $u(t)\in\IC^m$ is the input.
Assume that $(A,B)$ is controllable, $\Pi$ is Hermitian, and
$\bfOme$ has a nonempty interior. Then the following statements
are equivalent.
\begin{itemize}
\item[(i)] The frequency domain inequality (\ref{fdic})
holds for $\omega\in\bfOme$.
\item[(ii)] The time domain inequality
\begin{equation} \mylabel{tdic}
\int_0^\infty
\begin{mat}{c} x \\ u \end{mat}^*\Pi
\begin{mat}{c} x \\ u \end{mat}\dt\leq0
\end{equation}
holds for all solutions of (\ref{sysc}) with $u\in\Ltwo$ such that
$x(0)=0,~x\in\Ltwo$ and
\begin{equation} \mylabel{iqcc}
\he~\int_0^\infty (\varpi_1x+\j\dot x)(\varpi_2x+\j\dot
x)^*\dt\leq0.
\end{equation}
\end{itemize}
\end{theorem}
\medskip

Note that the corresponding result of \cite{IHF05} was obtained
under additional condition of asymptotic stability for the system
(\ref{sysc}) which is not required in the current statement.

 In Theorem~\ref{thm:tdic}, a general
frequency interval $\bfOme$ is considered for the FDI, and this
has translated to the input constraint described by (\ref{iqcc}).
Though the physical meaning of this constraint may be not clear in
 general, it becomes clear for the following  special case.

\begin{coro} \mylabel{coro:tdic}
Let real matrices $A$, $B$, $\Pi$, and a positive scalar $\varpi$
be given. Suppose $\Pi$ is symmetric and consider the system
(\ref{sysc}) where $x(t)\in\IR^n$ is the state and $u(t)\in\IR^m$
is the input. Assume that $(A,B)$ is controllable. Then the
following statements are equivalent.
\begin{itemize}
\item[(i)] The frequency domain inequality (\ref{fdic}) holds for
all $\omega$ such that $|\omega|\leq\varpi$. \item[(ii)] The time
domain inequality (\ref{tdic}) holds for all $u\in\Ltwo$ such that
$x\in\Ltwo$ and
\begin{equation} \mylabel{iqccl}
\int_0^\infty \dot x\dot x\t\dt\leq
\varpi^2
\int_0^\infty x x\t\dt.
\end{equation}
\end{itemize}
Moreover, the above two statements are equivalent when the two
inequalities ``$\leq$'' are replaced by ``$\geq$.''
\end{coro}

Loosely speaking, the first part of Corollary~\ref{coro:tdic}
states that the FDI in the low finite frequency range means that
the system possesses the property (\ref{tdic}) for the input
signals $u$ that  drive the states not too fast (slowly). The
bound on the ``slowness'' is given by $\varpi$ in the sense of
(\ref{iqccl}). The second part of Corollary~\ref{coro:tdic} makes
a similar statement for the FDI in the high frequency range. To
derive Corollary~\ref{coro:tdic} from Theorem~\ref{thm:tdic} one
needs to put $\omega_1=-\omega_2=\varpi$.

%It should be noted that condition (\ref{iqccl}) is a matrix-valued
%integral quadratic constraint, and is coordinate free, i.e. if
%(\ref{iqccl}) holds for some minimal realization of the
%system with states $x$, then it holds for all other minimal realizations
%with states $Tx$ where $T$ is an arbitrary nonsingular matrix.

{\it Proof of Theorem  ~\ref{thm:tdic}.}  In view of Theorem
\ref{thm:ckyp}  it is sufficient to prove equivalence of (ii) to
the condition (ii) of Theorem ~\ref{thm:tdic} (solvability of
matrix inequality  (\ref{lmic})). The result follows from Theorem
\ref{theorem:2} of the previous section with $m=2$ (one matrix
constraint). Denote
%\begin{array}
\begin{align} \mylabel{FzA}
z=&~\begin{mat}{c} x \\ u \end{mat}, ~F(z)=-\int_0^\infty
\begin{mat}{c} x \\ u \end{mat}^\dagger\Pi
\begin{mat}{c} x \\ u \end{mat}\dt, \nonumber
\\
G_1(z)&=-\he~\int_0^\infty (\varpi_1x+\j\dot x)(\varpi_2x+\j\dot
x)^\dagger\dt.
\end{align}
%\end{array}

Obviously, TDI (\ref{tdic}),(\ref{iqcc}) correspond to the
statement (A)  of the conic S-procedure. At the same time the
statement (B) means existence of a $n\times n$-matrix $\tau^*$
from the dual cone $K^*$ to the cone of positive semidefinite
matrices satisfying inequality $F(z)-<\tau^*, G_1(z)>\ge 0
~~\forall z$ or
\begin{equation} \mylabel{FzB}
%z=~\begin{mat}{c} x \\ u \end{mat},
\he\int_0^\infty \left(-
\begin{mat}{c} x \\ u \end{mat}^\dagger\Pi
\begin{mat}{c} x \\ u \end{mat}-(\varpi_1x+\j\dot x)^*\tau(\varpi_2x+\j\dot
x)\right)\dt\ge 0.
\end{equation}
Replacement of $\tau$ by $Q$ and substitution of $\dot x$ from
(\ref{sysc}) transforms (\ref{FzB}) into LMI (\ref{lmic}). To
verify regularity condition of Theorem \ref{theorem:2} take
$x_0=(\bar x,\bar u)$, where $$\bar x(t)=-(A+\mu I)^{-1}B\exp(-\mu
t),~\bar u(t)= \exp(-\mu t), ~\mu>0$$ and $\mu\in\inte \bfOme$.
 Application of Theorem  \ref{theorem:2}  ends the proof.

{\it Remark}. Similar results hold for discrete-time case.

\section{Conclusions}

The property of the system defined by the item (ii) of  Theorem
~\ref{thm:tdic} and the Corollary can be called {\it constrained
dissipativity} or {\it restricted dissipativity}. It is weaker
than standard passivity or dissipativity conditions and may better
reflect specifications for real systems. At the same time the
property of ``slowness" described by the inequality (\ref{iqccl})
leaves enough flexibility to be useful for robustness analysis of
systems.

 The results of the
paper shed new light on the intimate interrelations between
S-procedure and KYP-lemma. They allow to extend classical
S-procedure tool to allow for analysis and design of robust
systems with matrix inequalities constraints.

\end{document}